# Black hole interiors of homogeneous holographic solids under shear strain

### v2

@@ 1 Introduction
Since the Einstein equation was introduced in general relativity, the theoretical understanding of black hole exteriors has become well-established. In recent years, observational breakthroughs, including gravitational-wave detection LIGOScientific:2016aoc and direct imaging of black hole shadows EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ths , have further enriched our empirical knowledge of these regions. However, the physics of black hole interiors remains far less understood, primarily due to the presence of a curvature singularity enclosed by the event horizon, where conventional physical laws cease to hold. The inevitability of such singularities within general relativity was rigorously established by Penrose and Hawking Penrose:1964wq ; Hawking:1970zqf . In contrast to the spacelike singularity of the Schwarzschild solution, the singularity of a Reissner-Nordström (RN) black hole or a Kerr black hole is timelike owing to the presence of a Cauchy horizon. Moreover, the strong cosmic censorship (SCC) conjecture Isenberg:2015rqa asserts that any Cauchy horizon must be dynamically unstable to preserve the causal integrity of the spacetime Ringstrom:2015jza .
@@
A seminal contribution to the understanding of black hole interiors is the Belinski-Khalatnikov-Lifshitz (BKL) conjecture Lifshitz:1963ps ; Belinsky:1970ew ; Belinski:1973zz , which describes the dynamical behavior of spacetime near a spacelike singularity. According to this, the geometry undergoes chaotic, oscillatory dynamics characterized by an infinite sequence of Kasner epochs—periods of anisotropic expansion or contraction governed by distinct Kasner exponents. For a large number of models, the alternation between these epochs can be elegantly described using a cosmological billiard picture Damour:2002et ; Henneaux:2007ej , wherein the system’s evolution is mapped to a billiard ball reflecting off the walls of a multi-dimensional potential. This analogy provides a powerful tool for analyzing the chaotic nature of the BKL singularity and has profound implications for both the physics of black hole interiors and the early-universe cosmology.
@@
Another motivation for studying black hole interiors stems from the holographic duality Maldacena:1997re , which remarkably links weakly-coupled gravities to strongly coupled quantum field theories, has emerged as a transformative framework across disciplines, while providing novel approaches to study black hole interiors. Some dual field theory quantities, in particular, correlation functions Fidkowski:2003nf ; Festuccia:2005pi , entanglement entropy  Hartman:2013qma and quantum complexity Stanford:2014jda ; Brown:2015bva serve as complementary non-perturbative probes of interior dynamics.
@@
Recent studies have revealed that coupling matter fields to AdS-Schwarzschild Frenkel:2020ysx ; Hartnoll:2020rwq or RN black holes Hartnoll:2020fhc can dramatically modify their interior dynamics, particularly through the elimination of Cauchy horizons Cai:2020wrp ; An:2021plu ; Hale:2025ezt . The resulting dynamical evolution typically progresses through three characteristic phases: collapse of Einstein-Rosen (ER) bridge, Josephson oscillations and Kasner behavior near the singularity Hartnoll:2020fhc - with the latter potentially exhibiting Kasner inversions, transitions and reflections Cai:2023igv ; Cai:2024ltu . These phenomena have been verified in a broad class of systems within the framework of holography, including holographic superconductors Hartnoll:2020fhc ; Sword:2022oyg ; Cai:2021obq ; Cai:2024ltu ; Zhang:2025tsa ; Zhang:2025hkb , holographic semimetals Gao:2023zbd , holographic axion models Mansoori:2021wxf ; Mirjalali:2022wrg ; Prihadi:2025rwg ; Sword:2021pfm as well as the modified gravity with higher-derivative corrections Devecioglu:2021xug ; Devecioglu:2023hmn ; Grandi:2021ajl ; Caceres:2024edr ; Bueno:2024fzg . A common conclusion from these investigations is the necessity of coupling additional matter fields to the original gravitational theory to produce richer interior structures.
@@
Building on these developments, we examine black hole interiors within anisotropic holographic axion models Baggioli:2021xuv , with two crucial distinctions from previous work Mansoori:2021wxf ; Mirjalali:2022wrg ; Prihadi:2025rwg ; Sword:2021pfm : (1) our construction utilizes only the axion fields without additional (charged) scalar hair; and (2) the bulk axion fields realize the anisotropy on the boundary in a specific manner that allows finite shear deformations to be introduced in the holographic description [^1] Previous studies focused on isotropic cases in the presence of axions, for which a proof for no inner-horizon is absent.. This model has been applied to study nonlinear elasticity and mechanical failure of amorphous solids Baggioli:2020qdg ; Pan:2021cux ; Ji:2022ovs ; Baggioli:2023dfj , and more recently, it uncovers an unexpected connection between non-linear elasticity and quantum complexity Xu:2025uzo . However, in this work, we focus instead on its implications for black hole interiors - particularly how shear deformations, crucial to analyzing the elasticity of boundary systems, manifest in interior spacetime dynamics.
@@ 2 Holographic model
$\mathcal{I}^{IJ}\equiv\partial_{\mu}\phi^{I}\partial^{\mu}\phi^{J},\ \ X\equiv\frac{1}{2}\text{Tr}\left[\mathcal{I}^{IJ}\right],\ \ Z\equiv\det[\mathcal{I}^{IJ}],\ \ I,J=\{x,y\},$ (2)
@@
$\bar{\phi}^{I}=M^{I}_{i}x^{i},\quad i=x,y,$ (3)
where $M^{I}_{i}$ is a $2\times 2$ symmetric matrix. This class of models is nothing but the Stückelburg formalism of the Lorentz-violating massive gravity dual to broken translations on boundary Vegh:2013sk .
From the boundary perspective, the profile of axions $\phi^{I}\sim x^{i}$ cannot determine whether the breaking of translations is spontaneous or explicit. Instead, this should be fixed by the specific choice of the potential $V$. Sticking to the standard quantization, if $V$ decays much faster than $u^{5}$ near the boundary ($u=0$), there exist gapless phonon excitations with the speeds correlated with the elastic response of the system which corresponds to the spontaneous symmetry breaking (SSB) of translations Alberte:2017oqx . [^4] One can have a well-defined elastic response for potentials that decay at the boundary as $u^{3}$ or faster. In this case, the gravity system is dual to a homogeneous solid on boundary. On the contrary, if $V$ decays slower than $u^{5}$ near the boundary, the symmetry breaking is explicit which leads to momentum relaxation Andrade:2013gsa . The most part of this work will focus on the former case. However, an extended discussion about the later case will be attached at the end.
@@
where $\bar{X}=u^{2}\cosh(\Omega-h(u))$ and $\bar{Z}=u^{4}$. The Hawking temperature and the entropy density can be obtained from
@@ 3 Shear deformations remove the Cauchy horizon
Suppose that there were two horizons: the event horizon at $u_{h}$ and an inner horizon at $u_{i}$, such that
@@
Between the two, the blackening factor satisfies $f(u)<0$. From Eq. (10), we derive the following equation:
@@
$0$ $=$ $\int^{u_{i}}_{u_{h}}\left[\frac{e^{-\chi(u)/2}f(u)h^{\prime}(u)}{u^{2}\sinh[\Omega-h(u)]}\right]^{\prime}du,$ (16)
@@
The first equality follows from the condition $f(u_{h})=f(u_{i})=0$. Note that the product of the two hyperbolic functions in the last line is non-negative, i.e.,
$\coth[\Omega-h(u)]\text{csch}[\Omega-h(u)]\geq 0\,.$ (17)
Therefore, for the potential satisfying $V_{X}\geq 0$ as required by theoretical consistency (see Appendix A), the integrand in Eq. (16) remains non-positive throughout the integration domain. According to Eq. (12), the existence of two distinct horizons requires the isotropic limit with no shear deformation. This demonstrates that in the general case, anisotropic shear strain necessarily eliminates the inner horizon.
@@ 4.1 Collapse of Einstein-Rosen bridge under small shear deformations
Figure 3: (a): The behavior of the metric component $g_{tt}$ near the would-be inner horizon under different values of $\Omega$. The black dashed curves show semi-analytical results obtained from Eq. (22). (b): The behavior of the function $h^{\prime}(u)$ with respect to the radial coordinate $u/u_{h}$ inside the black hole. According to Eq. (25), the maximum value of $h^{\prime}$ near the would-be inner horizon corresponds to the ratio $c_{2}/c_{1}$. We have chosen $V(X,Z)=X^{2}\sqrt{Z}$ at $T/m=0.1$.
@@
$2\chi^{\prime}$ $=$ $u_{i}h^{\prime 2},$ (18)
$2u_{i}f^{\prime}$ $=$ $-6+u_{i}f\chi^{\prime},$ (19)
@@
To solve the equations, we begin by integrating Eq. (18) and writing $h^{\prime}=c_{1}\sqrt{-12}e^{\chi/2}/f$, where $c_{1}$ is a constant. The solution can be expressed in terms of the metric component $g_{tt}=-fe^{-\chi}/u_{i}^{2}$, which satisfies
@@
where $c_{2}$ and $c_{3}$ are integration constants (normalized for later convenience). In addition to the relation $f=-e^{\chi}g_{tt}u^{2}_{i}$, we obtain
$h=-\frac{4c_{1}}{u_{i}c_{2}}\log(c_{4}g_{tt}),\ \ e^{-\chi}=-\frac{u^{4}_{i}}{12c^{2}_{1}}h^{\prime 2}g^{2}_{tt}\,$, (23)
@@
Figure 4: The ratio $c_{2}/c_{1}$ as a function of $\Omega$, extracted via Eq. (25) near the would-be inner horizon. As $\Omega$ decreases, $c_{2}/c_{1}$ increases, resulting in a more dramatic crossover. Red dots and the black line represent the numerical data and the analytical approximation $c_{2}/c_{1}\simeq 32.8755/\Omega$, respectively, for the potential $V=X^{2}\sqrt{Z}$ at $T/m=0.1$.
The first equation in (23) reveals that an increasing $c_{2}/c_{1}$ ratio corresponds to decreasing shear deformation. In fact, the ratio exhibits an inverse proportionality to $\Omega$ in the small shear deformation regime, as shown in Fig. 4. This implies that, under small shear deformation, the ratio $(c_{2}/c_{1})^{2}\delta u$ remains significantly large in Eq. (24) even for small $\delta u$ near the would-be inner horizon. Consequently, the metric component $g_{tt}$ undergoes a drastic change within an extremely narrow range of the $u$-coordinate. Specifically, the second term on the left-hand side of Eq. (22) dominates for $u<u_{i}$ since $c^{2}_{1}/c^{2}_{2}\ll g_{tt}$. The first term becomes dominant when $c^{2}_{1}/c^{2}_{2}\gg g_{tt}$ for $u>u_{i}$. We summarize the abrupt changes in spacetime geometry during the rapid collapse of the ER bridge as follows:
$u<u_{i}$ $\rightarrow$ $u>u_{i}\,$,
$g_{tt}=\frac{u_{i}c^{2}_{2}}{4}|\delta u|$ $\rightarrow$ $g_{tt}=e^{-\frac{c^{2}_{2}u_{i}}{4c^{2}_{1}}\delta u}\,$, (24)
$h^{\prime}=\frac{c_{1}}{c_{2}|\delta u|}$ $\rightarrow$ $h^{\prime}=\frac{c_{2}}{c_{1}}\,$, (25)
where we have set the shift $c_{3}=0$ for clarity. The linear decay of $g_{tt}$ near the would-be inner horizon transitions abruptly to an exponential suppression, whereas $h^{\prime}(u)$, which would otherwise diverge at the horizon, saturates at a finite magnitude. As exhibited in Fig. 3, the numerical result shows excellent agreement with the aforementioned analytical analysis.

### v3

@@ 1 Introduction
Since the establishment of general relativity, understanding the nature of black holes has become a significant scientific question. In recent years, observational breakthroughs, including gravitational-wave detection LIGOScientific:2016aoc and direct imaging of black hole shadows EventHorizonTelescope:2019dse ; EventHorizonTelescope:2019ths , have further enriched our empirical knowledge of black hole exteriors. However, the physics of black hole interiors remains far less understood, primarily due to the presence of a curvature singularity enclosed by the event horizon, where conventional physical laws cease to hold. The inevitability of such singularities within general relativity was rigorously established by Penrose and Hawking Penrose:1964wq ; Hawking:1970zqf . In contrast to the spacelike singularity of the Schwarzschild solution, the singularity of a Reissner-Nordström (RN) black hole or a Kerr black hole is timelike owing to the presence of a Cauchy horizon. Moreover, the strong cosmic censorship (SCC) conjecture Isenberg:2015rqa asserts that any Cauchy horizon must be dynamically unstable to preserve the causal integrity of the spacetime Ringstrom:2015jza .
@@
A seminal contribution to the understanding of black hole interiors is the Belinski-Khalatnikov-Lifshitz (BKL) conjecture Lifshitz:1963ps ; Belinsky:1970ew ; Belinski:1973zz which describes the dynamical behavior of spacetime near a spacelike singularity. According to this, the geometry undergoes chaotic, oscillatory dynamics characterized by an infinite sequence of Kasner epochs—periods of anisotropic expansion or contraction governed by distinct Kasner exponents. For a large number of models, the alternation between these epochs can be elegantly described using a cosmological billiard picture Damour:2002et ; Henneaux:2007ej , wherein the system’s evolution is mapped to a billiard ball reflecting off the walls of a multi-dimensional potential. This analogy provides a powerful tool for analyzing the chaotic nature of the BKL singularity and has profound implications for both the physics of black hole interiors and the early-universe cosmology.
@@
The study of black hole interiors is further motivated by holographic duality—a profound theoretical link between weakly coupled gravitational systems and strongly coupled quantum field theories. As a cross‑disciplinary framework, it also provides novel pathways for probing the interior structure of black holes. Some dual field theory quantities, in particular, correlation functions Fidkowski:2003nf ; Festuccia:2005pi , entanglement entropy  Hartman:2013qma and quantum complexity Stanford:2014jda ; Brown:2015bva serve as complementary non-perturbative probes of interior dynamics.
@@
Recent studies have revealed that coupling matter fields to AdS-Schwarzschild Frenkel:2020ysx ; Hartnoll:2020rwq or RN black holes Hartnoll:2020fhc can dramatically modify their interior dynamics, particularly through the elimination of Cauchy horizons Cai:2020wrp ; An:2021plu ; Hale:2025ezt . The resulting dynamical evolution typically progresses through three characteristic phases: collapse of Einstein-Rosen (ER) bridge, Josephson oscillations and Kasner behavior near the singularity Hartnoll:2020fhc - with the latter potentially exhibiting Kasner inversions, transitions and reflections Cai:2023igv ; Cai:2024ltu . These phenomena have been verified in a broad class of systems within the framework of holography, including holographic superconductors Hartnoll:2020fhc ; Sword:2022oyg ; Cai:2021obq ; Cai:2024ltu ; Zhang:2025tsa ; Zhang:2025hkb , holographic metals Gao:2023zbd ; Carballo:2024hem , holographic axion models Mansoori:2021wxf ; Mirjalali:2022wrg ; Prihadi:2025rwg ; Sword:2021pfm ; Arean:2024pzo as well as the modified gravity with higher-derivative corrections Devecioglu:2021xug ; Devecioglu:2023hmn ; Grandi:2021ajl ; Caceres:2024edr ; Bueno:2024fzg . A common conclusion from these investigations is the necessity of coupling additional matter fields to the original gravitational theory to produce richer interior structures.
@@
Building on these developments, we examine black hole interiors within anisotropic holographic axion models Baggioli:2021xuv , with two crucial distinctions from previous work Mansoori:2021wxf ; Mirjalali:2022wrg ; Prihadi:2025rwg ; Sword:2021pfm ; Arean:2024pzo : (1) our construction utilizes only the axion fields without additional (charged) scalar hair; and (2) the bulk axion fields realize the anisotropy on the boundary in a specific manner that allows finite shear deformations to be introduced in the holographic description [^1] Previous studies focused on isotropic cases in the presence of axions, for which a proof for no inner-horizon is absent Mirjalali:2022wrg ; Prihadi:2025rwg ; Sword:2021pfm .. This model has been applied to study nonlinear elasticity and mechanical failure of amorphous solids Baggioli:2020qdg ; Pan:2021cux ; Ji:2022ovs ; Baggioli:2023dfj , and more recently, it uncovers an unexpected connection between non-linear elasticity and quantum complexity Xu:2025uzo . However, in this work, we focus instead on its implications for black hole interiors - particularly how shear deformations, crucial to analyzing the elasticity of boundary systems, manifest in interior spacetime dynamics.
@@ 2 Holographic model
$\mathcal{I}^{IJ}\equiv\partial_{\mu}\phi^{I}\partial^{\mu}\phi^{J},\ \ X\equiv\frac{1}{2}\text{Tr}\left[\mathcal{I}^{IJ}\right],\ \ Z\equiv\det[\mathcal{I}^{IJ}],\ \ I,J=\{1,2\},$ (2)
@@
$\bar{\phi}^{I}={M^{I}}_{i}x^{i},\quad i=\{1,2\},$ (3)
where ${M^{I}}_{i}$ is a $2\times 2$ symmetric matrix, the notations $x^{1}\equiv x$ and $x^{2}\equiv y$ have been adopted. This class of models is nothing but the Stückelburg formalism of the Lorentz-violating massive gravity dual to broken translations on boundary Vegh:2013sk .
From the boundary perspective, the profile of axions $\phi^{I}\sim x^{i}$ cannot determine whether the breaking of translations is spontaneous or explicit. Instead, this should be fixed by the specific choice of the potential $V$. Sticking to the standard quantization, if $V$ decays much faster than $u^{5}$ near the boundary ($u=0$), there exist gapless phonon excitations with the speeds correlated with the elastic response of the system which corresponds to the spontaneous symmetry breaking (SSB) of translations Alberte:2017oqx . [^4] One can have a well-defined elastic response for potentials that decay at the boundary as $u^{3}$ or faster. In this case, the gravity system is dual to a homogeneous solid on boundary. On the contrary, if $V$ decays slower than $u^{5}$ near the boundary, the symmetry breaking is explicit which leads to momentum relaxation Andrade:2013gsa . For the relationship between the mechanism of translational symmetry breaking and the near-boundary behavior of $V$, one may refer to Appendix A for further details. The most part of this work will focus on the former case. However, an extended discussion about the later case will be attached at the end.
@@
where $\bar{X}=u^{2}\cosh[\Omega-h(u)]$ and $\bar{Z}=u^{4}$. The Hawking temperature and the entropy density can be obtained from
@@ 3 Shear deformations remove the Cauchy horizon
Suppose that there exist two horizons: an event horizon located at $u_{h}$ and an inner horizon at $u_{i}$, such that
@@
Between these two horizons, the blackening factor satisfies $f(u)<0$. For the black hole under shear strain (with nonzero $h$ and $\Omega$), the following equation can be derived from Eq. (10):
@@
$0$ $=$ $\int^{u_{i}}_{u_{h}}\left[\frac{e^{-\chi(u)/2}f(u)h^{\prime}(u)}{u^{2}\sinh[\Omega-h(u)]}\right]^{\prime}du$ (16)
@@
The first equality follows from the condition $f(u_{h})=f(u_{i})=0$. Note that the product of the two hyperbolic functions in the last line is always positive, i.e.,
$\coth[\Omega-h(u)]\text{csch}[\Omega-h(u)]>0\,.$ (17)
Therefore, for the potential satisfying $V_{X}\geq 0$ as required by theoretical consistency (see Appendix A), the integrand in Eq. (16) remains negative throughout the integration domain. According to Eq. (12), the existence of two distinct horizons requires the isotropic limit with no shear deformation. This demonstrates that in the general case, anisotropic shear strain necessarily eliminates the inner horizon.
@@ 4.1 Collapse of Einstein-Rosen bridge under small shear deformations
Figure 3: (a): The behavior of the metric component $g_{tt}$ near the would-be inner horizon under different values of $\Omega$. The black dashed curves show semi-analytical results obtained from Eq. (22). (b): The behavior of the function $h^{\prime}(u)$ with respect to the radial coordinate $u/u_{h}$ inside the black hole. According to Table. 1, the maximum value of $h^{\prime}$ near the would-be inner horizon $u_{i}$ corresponds to the ratio $c_{2}/c_{1}$. We have chosen $V(X,Z)=X^{2}\sqrt{Z}$ at $T/m=0.1$.
@@
$2\chi^{\prime}$ $=$ $u_{i}(h^{\prime})^{2},$ (18)
$4u_{i}f^{\prime}$ $=$ $u^{2}_{i}f(h^{\prime})^{2}+4m^{2}V(u^{2}_{i},u^{4}_{i})-12,$ (19)
@@
To solve the equations, we begin by integrating Eq. (20) and writing
$h^{\prime}=-c_{1}\left[4m^{2}V(u^{2}_{i},u^{4}_{i})-12\right]^{1/2}e^{\chi/2}/f$, where $c_{1}$ is a constant. The solution can be expressed in terms of the metric component $g_{tt}=-fe^{-\chi}/u_{i}^{2}$, which satisfies
@@
where $c_{2}$ and $c_{3}$ are integration constants (normalized for later convenience). In addition to relations $f=-e^{\chi}g_{tt}u^{2}_{i}$ and $c^{2}_{1}g^{\prime}_{tt}/g_{tt}=-u_{i}(c^{2}_{1}+g_{tt})(h^{\prime})^{2}/4$, we obtain
$h=-\frac{4c_{1}}{u_{i}c_{2}}\log(c_{4}g_{tt}),\ \ e^{-\chi}=\frac{u^{4}_{i}}{c^{2}_{1}\left[4m^{2}V(u^{2}_{i},u^{4}_{i})-12\right]}(h^{\prime})^{2}g^{2}_{tt}\,$, (23)
@@
Figure 4: The ratio $c_{2}/c_{1}$ as a function of $\Omega$, extracted via $h^{\prime}=c_{2}/c_{1}$ in Table 1 near the would-be inner horizon. As $\Omega$ decreases, $c_{2}/c_{1}$ increases, resulting in a more dramatic crossover. Red dots and the black line represent the numerical data and the analytical approximation $c_{2}/c_{1}\simeq 32.8755/\Omega$, respectively, for the potential $V=X^{2}\sqrt{Z}$ at $T/m=0.1$.
The first equation in Eq. (23) reveals that an increasing $c_{2}/c_{1}$ ratio corresponds to decreasing shear deformation. In fact, ratio $c_{2}/c_{1}$ exhibits an inverse proportionality to $\Omega$ in the small shear deformation regime, as shown in Fig. 4. For a sufficiently small shear deformation ($\Omega\ll 1$), in the region outside (but in the vicinity of) the would-be inner horizon (i.e., $\delta u<0$), it is evident that $g_{tt}\gg c^{2}_{1}/c^{2}_{2}$. Therefore, according to Eq. (22) and the first equation of Eq. (23), we find $g_{tt}=\frac{u_{i}c^{2}_{2}}{4}|\delta u|$ and $h^{\prime}=\frac{4c_{1}}{u_{i}c_{2}}\frac{1}{|\delta u|}$. Similarly, in inside the would-be inner horizon (i.e. $\delta u>0$), where $g_{tt}\ll c^{2}_{1}/c^{2}_{2}$, we obtain $g_{tt}=\text{Exp}\left[-\frac{u_{i}c^{2}_{2}}{4c^{2}_{1}}\delta u\right]$ and $h^{\prime}=c_{2}/c_{1}$. We summarize the abrupt changes in spacetime geometry during the rapid collapse of the ER bridge in Table 1. The linear decay of $g_{tt}$ near the would-be inner horizon transitions abruptly to an exponential suppression, while $h^{\prime}(u)$ saturates abruptly from a very small value to a large, finite magnitude under small shear deformation.
[…]
$g_{tt}=\frac{u_{i}c^{2}_{2}}{4}|\delta u|$     $g_{tt}=e^{-\frac{u_{i}c^{2}_{2}}{4c^{2}_{1}}\delta u}$
$h^{\prime}=\frac{4c_{1}}{u_{i}c_{2}}\frac{1}{|\delta u|}$ $h^{\prime}=\frac{c_{2}}{c_{1}}$
Table 1: The metric component $g_{tt}$ and $h^{\prime}$ near the would-be inner horizon $u_{i}$. For clarity, the shift parameter $c_{3}$ in Eq. (22) has been set to zero.
Fig. 4 shows the ratio $c_{2}/c_{1}$ as a function of $\Omega$, extracted from the relation $h^{\prime}=c_{2}/c_{1}$ near the would-be inner horizon. The ratio exhibits an inverse proportionality to $\Omega$ in the small-deformation regime. This implies that under small shear deformation, the quantity $(c_{2}/c_{1})^{2}\delta u$ remains large even for small $\delta u$ near the horizon. Consequently, both the metric component $g_{tt}$ and the anisotropic derivative $h^{\prime}$ undergo an abrupt change within an extremely narrow interval of the $u$-coordinate, as demonstrated by the numerical results in Fig. 3 and the analytic results in Table 1. As shown in Fig. 3(a), the numerical date exhibit excellent agreement with the preceding analytical analysis.